$SPAD/src/input axiom.input

The Axiom Team December 3, 2016

Abstract

Contents

2

1 License

— license —

--Copyright The Numerical Algorithms Group Limited 1994.

———-

— * —

\getchunk{license}
-- This file is read on start-up by the AXIOM system.  It is
-- most commonly used to customize the user's environment via the
-- )set system commands.

-- do not enter lisp break on error
)set break nobreak
-- use the simple character set
)set output char plain
-- everyone uses this
)set userlevel development
-- make the highlighting characters go away
)set mess high off
)set message prompt frame
-- read the user's synonym file for system commands, if it exists
)read "mysyns.input"  )ifthere )quiet

———-

References^{[1] nothing}

4